Present: Paki, Ivan, Stefano, Alex O., Yvonne, Alice, Peter

Apologies: Marco, Jo



Clean-rooms rearrangement: this week we were meant to move the Diamond stuff from the right room to the middle one, but most of the people aren't available. It is decided to postopone it to Thursday 29th from 2pm to 6pm. Volunteers: Alex O., Paki, Alice, Stefano and Ivan.

The new probe station will replace the old one in the right clean-room, therefore we need to decide what to to with the old one. 



Roundtable:



Paki is progressing with the purchase of a new probe station. Due to "fear of Brexit", the company wants a purchase order by the end of this week and in return they have offered us quite a good deal. Paki is unsured though if all documents can be prepared in time. Currently it is also not clear how the cooling system (for the -40C chuck) will be done: the company proposed compressed air, but our dryer cannot cope with the requested flow (550 l/min). A new dryer costs 6k £.

Paki is also busy with her application for a fellowship, deadline on this Friday. 



Ivan is working on the 2-photon absorption-TCT at the PSI building. Main issue is that the power of laser, currently set to create columns inside the diamond, is too high and has to be reduced. Ivan went to Ljubjana for a RD42 meeting and gave a talk on his work about 3D diamond processing.



Stefano reports that the work on the VELO Upgrade is progressing and they are currently working on their 2nd electrical prototype. Mike has successfully wire-bonded a whole side of a mechanical module in one go and he's now about to move to the electrical. The availability of the final parts for module production is still an issue and it will likely postpone the Site Qualification Review to February.



Peter will give atalk on light detection with TimePix3 at the next Medipix collaboration meeting. He will also go to the next RD50 meeting but he's not meant to give a talk there. He believes Gedas will present the results on the VELO radiation damage. Peter's also getting involved in the R&D of LGAD detectors for the next VELO Upgrade. 



Alex has been in Ljubjana as well. He's also working on the BCM for the Atlas Upgrade.



Yvonne reports that the two MPhys students working with Paki on the thermal qualification of the Atlas rings are doing very well and seem to quite enjoy the Lab. They are now working on Niko's software for the thermal measurements. Paki will try to get them to present their work at the next meeting.



Alice is currently debugging Giulio's software on simulation.